Show that the following series converges: (4)
(2n 3)(2n 5)
1
n 1 + + ∑

=
1
Expert's answer
2019-10-16T09:25:43-0400

Proof:

We need to show that the series converges.

Let "\displaystyle\sum_{i=1}^n""\frac {1} {(2n+3) (2n+5)} =" "\displaystyle\sum_{i=1}^n" ui Here un ="\frac {1} {(2n+3) (2n+5)}"



Let us consider the series "\displaystyle\sum_{i=1}^n" vi here vn ="\frac {1} {n^2}"



Now, "\lim\limits_{n->\infin}""\frac {} { }""\frac {u_n} {v_n}" = "\lim\limits_{n->\infin}" "\frac {\frac {1} {[(2n+3) (2n+5)]}} {\frac {1} {(n^2)}}" = "\lim\limits_{n->\infin}" "\frac {n^2} {(2n+3) (2n+5)}"



= "\lim\limits_{n->\infin}""\frac {1} {(2+\frac {3} {n} ) (2+\frac {5} {n})}" = "\frac {1} {(2+0) (2+ 0) }" ( Here "\frac {3} {\infin } = 0" and "\frac {5} {\infin } = 0" )


= "\frac {1} {4} < 1"


"\lim\limits_{n->\infin}""\frac {u_n} {v_n} < 1" , then the given series converges.

Answer: the series converges.